Line A: y = x − 2 Line B: y = 3x + 4 Which of

Line A: y = x − 2Line B: y = 3x + 4

Which of the following shows the solution to the system of equations and explains why?
Answer #1

y = x -2 ......(1)

y= 3x + 4 ......(2)


(2) - (1)


0 = 3x + 4 -x + 2

=> 2x = -6

=> x = -3

y = x-2 = -5


answered by: dency